OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 A. slope = 5 over 3; y-intercept = (0, 2) B. slope = -5; y-intercept = (0, -6) C. slope = negative 3 over 5; y-intercept = (0, 2) D. slope = 5 over 3; y-intercept = (0, -6)

OpenStudy (anonymous):

At fist take it to the y-intercept form y=mx+c

OpenStudy (anonymous):

Can u?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

5x - 3y = -6 -3y=-6-5x 3y=5x+6 y=5/3 x+2

OpenStudy (anonymous):

Now, slope = m and y-intercept=c

OpenStudy (anonymous):

ohh ok so the slope would be 5/3 and the y intercept would me 0,2?

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

omg thank u soo much!! (: